OpenStudy (nuccioreggie):

An equation is shown below: 5(2x – 3) = 5 Part A: How many solutions does this equation have? (4 points) Part B: What are the solutions to this equation? Show your work. (6 points)

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

5(2x-3)=5 10x-15=5 (distribute) 10x=20 (add 15 to both sides) x=2 (divide both sides by 10) There is only one solution, and it is x=2